OpenStudy (anonymous):

Can someone please help me with this question?Find the difference of (−5x2 − x3 − 3) − (9x3 + 3x − 6x2).

OpenStudy (anonymous):

is it -5x^2? or (-5)(2)?

OpenStudy (anonymous):

if you required simple difference its -10x^3+x^2-3x-3

OpenStudy (anonymous):

\[-5x^2 - x^3 -3 - 9x^3 -3x+6x^2 = -(9x^3 + x^3) + (6x^2-5x^2) - 3x -3\] Add them and solve.

OpenStudy (anonymous):

(-5x^2-x^3-3)-1(9x^3+3x-6x^2) = -5x^2-x^3-3-9x^3-3x+6x^2 = (-9-1)x^3+ (-5+6)x^2-3x-3 =mary92